OpenStudy (anonymous):

1. Given the reaction: 2 A1 + 3 H2SO4 --> A12(SO4)3 + 3 H2 If an excess of sulfuric acid is combind with 10.0g of aluminum, and 0.150g of hydrogen is obtained, what is the percentage yield of the hydrogen? EXPLAIN ALL STEPS THIS NEEDS TO BE DONE BY TONIGHT SO I HOPE I CAN GET HELP!

OpenStudy (aaronq):

Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To interconvert between mass and moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. ---------------------------------------------------------------- to find the percent yield, use: percent yield\(=\dfrac{experimental~yield}{theoretical~yield}*100\%\) the experimental yield is given in the question. Next time ask chemistry questions in the chemistry section.